### v1

# Energy Efficient Exact and Approximate Systolic Array Architecture for Matrix Multiplication

Abstract
Deep Neural Networks (DNNs) require highly efficient matrix multiplication engines for complex computations. This paper presents a systolic array architecture incorporating novel exact and approximate processing elements (PEs), designed using energy-efficient positive partial product and negative partial product cells, termed as PPC and NPPC, respectively. The proposed 8-bit exact and approximate PE designs are employed in a 8x8 systolic array, they achieve energy savings of 22% and 32%, respectively, compared to the existing design. To demonstrate their effectiveness, the proposed PEs are integrated into a systolic array (SA) for Discrete Cosine Transform (DCT) computation, achieving high output quality with a PSNR of 38.21 dB. Furthermore, in an edge detection application using convolution, the approximate PE achieves a PSNR of 30.45 dB. These results highlight the potential of the proposed design to deliver significant energy efficiency while maintaining competitive output quality, making it well-suited for error-resilient image and vision processing applications.
Index Terms:
Approximate Computing, Systolic Array, Matrix Multiplication, Processing Element, Low Power Design, DCT, Edge Detection
I Introduction
Artificial Intelligence (AI) and Machine Learning (ML) have become integral to a wide range of applications, including computer vision, natural language processing, speech recognition, and autonomous systems [1, 2]. At the core of these workloads are deep neural networks (DNNs), which rely extensively on large-scale matrix multiplications to perform computations across multiple layers of interconnected neurons. Both the training and inference phases of DNNs involve billions of multiply–accumulate (MAC) operations, making matrix multiplication one of the most computationally and energy-intensive kernels in modern AI accelerators [3].
Systolic arrays (SAs) have emerged as an effective hardware architecture for accelerating matrix multiplications [4]. Recent advances in domain-specific hardware, such as Google’s Tensor Processing Unit (TPU), have demonstrated the ability of systolic arrays to deliver high throughput for DNN workloads. However, conventional systolic arrays typically employ exact arithmetic units, which incur significant energy consumption and large silicon area overheads. These limitations hinder their deployment in resource-constrained platforms, such as edge devices and Internet of Things (IoT) systems, where energy efficiency is of paramount importance. To mitigate these challenges, approximate computing has been widely explored as a design paradigm that trades exactness for improvements in energy efficiency and hardware complexity.
In this paper, we propose both exact and approximate processing element that supports both signed and unsigned operations for systolic array-based matrix multiplication. The proposed PE significantly reduces hardware complexity and energy consumption while maintaining acceptable computational accuracy. This makes it well-suited for error-tolerant applications such as image processing and machine learning, where efficiency is prioritized over exact precision. The major contributions of this paper are summarized as follows:
  • Proposed an energy-efficient approximate systolic array architecture for matrix multiplication, introducing novel exact and approximate Processing Elements (PEs).
  • Proposed approximate Positive Partial Product Cell (PPC)s and Negative Partial Product Cell (NPPC)s, these blocks achieve 46.8% and 38.7% of energy savings respectively as compared to the best existing design [5].
  • Proposed exact signed 8-bit PE achieves 20.26% of energy saving over the design [6], while the proposed approximate signed 8-bit PE achieves 13.11% of energy saving over the design [5].
  • Validation of the proposed architecture on two applications: DCT for image compression and convolution-based edge detection, results have found to achieve a PSNR of 45 dB for DCT, 30.45 dB for edge detection.
The rest of the paper is organized as follows. Section II explains the matrix multiplication using systolic array architecture. Section III presents the proposed processing elements for the matrix multiplication. Sections IV and V present the results and applications, respectively. The conclusions are presented in Section VI.
II Systolic Array Based Matrix Multiplication Architecture
Matrix multiplication plays a major role in accelerating the applications such as machine learning, signal processing, and neural networks. Several systolic array architectures for matrix multiplication are found in literature [7, 8, 9]. Recently, with applications in neural networks with larger data, GEMM-based matrix multiplication using a systolic array is highly appreciated for machine learning applications [10], [8].
Figure 1: Architecture of $3\times 3$ Systolic Array for Matrix Multiplication[7].
The latency of the conventional systolic array architecture, illustrated in Fig. 1, is given by 3N23N-2 clock cycles [11].
In a conventional systolic array as shown in Fig. 1, the processing element (PE) is the basic block for executing multiply-accumulate (MAC) operations [9]. Since a large number of PEs operate concurrently, it has a critical impact on the overall performance, energy efficiency, and area utilization. In the conventional design, a PE typically consists of a single multiplier followed by an adder to perform the multiply-accumulate (MAC) operation [9]. Alternatively, some studies optimize the PE by employing Positive Partial Product Cells (PPC) and Negative Partial Product Cells (NPPC) [6], to generate positive partial products and negative partial products, respectively, for enabling efficient signed multiplication as presented in Fig. 2. The partial products are then accumulated using full adders (FA) in the accumulation stage. Fig. 3 illustrates the exact PPC and NPPC blocks used to perform the bit-wise MAC operations.
Figure 2: Existing Design of Exact 4-Bit Signed PE [6]
Figure 3: Conventional approach of Exact (a) PPC and (b) NPPC
The next section presents the proposed processing elements for performing the matrix multiplication.
III Hardware Implementation of Proposed Processing Element (PE)
First, we present the exact PE architecture using the proposed optimized PPC and NPPC cells.
III-A Proposed Exact Processing Element
The proposed PE architecture eliminates the conventional separation between multiplication and accumulation by integrating them in to a single processing element to perform $a\times b+c$. This fused approach enables simultaneous reduction of both partial products and the accumulated sum, improving efficiency and reducing delay. Fig. 4 illustrates the exact 4-bit unsigned PE. In this architecture, PPC play a crucial role in performing the MAC operation. The proposed PPC for the PE are shown in Fig. 6 (a).
Figure 4: Proposed Exact 4-Bit Unsigned PE Design
Figure 5: Proposed Exact 4-Bit Signed PE Design
Figure 6: Proposed exact designs of (a) PPC and (b) NPPC
Figure 7: Proposed exact 8-bit signed PE Design
Fig. 5 and 7 illustrates the exact 4 bit and 8-bit signed PE architecture, respectively. The NPPC cells play a crucial role in handling signed operations, and the proposed cell is shown in Fig. 6 (b). For the 8-bit PE, a total of 14 NPPC cells and 50 PPC cells are required, whereas the existing work in [6] requires 15 number of additional full adders as well. In general, for an $N$-bit PE, the implementation requires $N^{2}-2N-2$ PPC cells and 2N22N-2 NPPC cells.
III-B Proposed Approximate Processing Element
In the proposed design, the PE is constructed using approximate PPC and NPPC blocks. Fig.  8 illustrates the architecture of the proposed approximate 4-bit signed PE, where the dotted blocks represent the approximate PPC and NPPC units. The proposed approximate PPC and NPPC are presented in Fig.  9, and their truth tables are provided in Table I.
Figure 8: Proposed Approximate 4-Bit Signed PE Design
The Boolean expression for the Sum output of the proposed approximate PPC cell is $S_{\text{out}}=\overline{\overline{(S_{\text{in}}+C_{\text{in}})}+(a_{i}\cdot b_{i})}$, while the corresponding carry output is $C_{\text{out}}=a_{i}\cdot b_{i}$. For the NPPC block, the Sum output is expressed as $S_{\text{out}}=\overline{(S_{\text{in}}+C_{\text{in})}\cdot\overline{(a_{i}\cdot b_{i})}}$, and the carry output is given by $C_{\text{out}}=(S_{\text{in}}+C_{\text{in})}\cdot\overline{(a_{i}\cdot b_{i})}$. The proposed approximate PPC introduces five erroneous outputs, out of a total of 16 cases, resulting in an error rate of $5/16$. The error occurs for the following combinations of inputs, $(a_{i},b_{i},S_{\text{in}},C_{\text{in}})$ are $(0,0,1,1)$, $(0,1,1,1)$, $(1,0,1,1)$, $(1,1,0,0)$, and $(1,1,1,1)$. The corresponding error probabilities ($P_{E}$) for these cases are $9/256$, $3/256$, $3/256$, $9/256$, and $1/256$, respectively. Therefore, the total error probability of the proposed PPC and NPPC blocks is $25/256$.
Figure 9: Proposed Approximate designs of (a) PPC and (b) NPPC
TABLE I: Truth Table for Partial Product Cell (PPC) and Negative Partial Product Cell (NPPC)
Inputs PPC: $a\cdot b+C_{in}+S_{in}$ NPPC: $\overline{a\cdot b}+C_{in}+S_{in}$
Exact Approx Exact Approx
$a_{i}$ $b_{i}$ $C_{in}$ $_{in} SS$ $C$ $S$ $C$ ED $S$ $C$ $S$ $C$ ED
0 0 0 0 0 0 0 0 0 1 0 1 0 0
0 0 0 1 1 0 1 0 0 0 1 0 1 0
0 0 1 0 1 0 1 0 0 0 1 0 1 0
0 0 1 1 0 1 1 0 -1 1 1 0 1 -1
0 1 0 0 0 0 0 0 0 1 0 1 0 0
0 1 0 1 1 0 1 0 0 0 1 0 1 0
0 1 1 0 1 0 1 0 0 0 1 0 1 0
0 1 1 1 0 1 1 0 -1 1 1 0 1 -1
1 0 0 0 0 0 0 0 0 1 0 1 0 0
1 0 0 1 1 0 1 0 0 0 1 0 1 0
1 0 1 0 1 0 1 0 0 0 1 0 1 0
1 0 1 1 0 1 1 0 -1 1 1 0 1 -1
1 1 0 0 1 0 0 1 +1 0 0 1 0 +1
1 1 0 1 0 1 0 1 0 1 0 1 0 0
1 1 1 0 0 1 0 1 0 1 0 1 0 0
1 1 1 1 1 1 0 1 -1 0 1 1 0 -1
IV Results and Discussion
This section presents hardware evaluation of the proposed and existing designs in terms of area, power, and delay using the Cadence Genus synthesis solution. In addition, error metrics are analyzed through a Python-based simulations.
IV-A Hardware Evaluation
The existing and proposed designs are designed using Verilog HDL, and then the designs are synthesized using the Cadence Genus Synthesis solution using a 90-nm UMC technology. The synthesized results were then evaluated based on three primary hardware metrics: critical path delay, power consumption, and area, as summarized in Table II for the proposed PPC and NPPC. The analysis shows an energy improvement of 6.4% for the exact design [6], and a significant 46.8% improvement for the proposed approximate PPC design compared to the best existing design [12].
TABLE II: Hardware Analysis of Proposed and Existing PPCs and NPPCs
Design PPC NPPC
Area Power Delay PDP Area Power Delay PDP
[…]
Exact [6] 25.81 1.03 262 269.86 24.92 0.99 238 235.62
Prop Ext 24.98 0.99 255 252.45 23.47 0.99 216 213.84
Design [6] 13.32 0.64 187 119.04 12.54 0.61 156 95.16
Design [12] 14.13 0.58 157 91.06 13.22 0.60 148 88.80
Prop Apx 10.19 0.44 110 48.4 9.40 0.37 147 54.39
Table III presents the analysis of the proposed and existing PEs for both unsigned and signed designs. The results indicate that the proposed exact PEs achieve superior efficiency compared to existing exact designs, with notable reductions in area, power, and delay, leading to overall improvements of up to 16%. Table III also includes PE designs without PPC and NPPC blocks [13, 14]. A significant improvement in the Power Area Delay Product (PADP) of 65.45% is observed when compared to the design in [13], highlighting that PPC and NPPC based PE designs are highly efficient in terms of hardware metrics. Furthermore, the proposed approximate PE achieve up to 23% improvement in the PADP compared to the best existing approximate design [5].
TABLE III: Hardware Analysis of Proposed and Existing Processing Elements
PE Bit Unsigned Signed
Design Width Area Power Delay PADP Area Power Delay PADP
[…]
Exact PPCs and NPPCs based Designs
Design 4 435.9 29.4 1.87 23.96 446.5 29.7 1.65 21.82
[6] 8 1718.5 181.3 3.92 1222.57 1708 183.4 3.71 1162.39
Design 4 432.8 30.4 1.76 23.13 445.3 31.7 1.55 21.88
[5] 8 1730.6 185.3 3.67 1175.71 1716 190.3 3.22 1050.21
Proposed 4 411 26.6 1.73 18.91 419 26.8 1.52 17.06
8 1659.2 180.7 3.65 1094.33 1620.3 170.6 3.18 879.02
Other Exact Designs
[13] 8 - - - - 1891 481 2.8 2546.7
[14] 8 - - - - 2366 574 2.5 3395.2
Approximate PPCs and NPPCs based Designs
Design [6] 4 416.30 24.1 1.56 15.64 435.9 29.6 1.69 21.78
8 1557.5 172.2 3.55 950.04 1546.3 216.0 3.51 1171.47
Design [12] 4 407.68 25.5 1.43 14.85 427.28 31.7 1.61 21.88
8 1476.2 164.1 3.21 777.51 1465.2 207.9 3.18 966.75
Design [5] 4 412.2 25.8 1.40 14.90 420.1 28.3 1.40 16.64
8 1012.1 145.5 3.01 442.91 975.5 177.2 2.50 431.93
Proposed 4 375.6 17.1 1.37 8.79 399.3 25.6 1.35 13.79
8 985.2 125.3 2.71 334.53 869.5 155.2 2.48 334.66
Figure 10: Area and PDP comparison of the exact design [6] and the proposed design across systolic array sizes.
Table IV summarizes the results for various systolic array (SA) sizes at 250MHz250~MHz. Fig. 10 shows up to 5.9% area savings and 12.8% energy savings of the proposed exact design over [6] across various sizes of systolic array. For the 8-bit configuration with a 16×16 SA, the proposed approximate design achieves a 55.9% reduction in PDP compared to the exact design [6] and a 10.5% improvement compared to the best exciting approximate design [5]. These results indicate that the proposed approximate designs achieve substantial gains in hardware efficiency, particularly for large systolic arrays, while maintaining competitive accuracy-performance tradeoffs. Fig. 11 compares the PADP across different systolic array sizes, where the proposed approximate design consistently outperforms existing designs [6], [5].
TABLE IV: Hardware Analysis of Different Designs for Various Systolic Array Sizes at 250 MHz
Bit Design 3×3 4×4 8×8 16×16
Width Area Power Delay PDP Area Power Delay PDP Area Power Delay PDP Area Power Delay PDP
(mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ)
4 Exact [6] 0.0062 3.98 1.65 6.57 0.0112 3.98 1.67 6.65 0.0465 17.2 1.88 32.34 0.1901 74.4 2.41 179.30
Proposed Exact 0.0060 3.90 1.63 6.35 0.0110 3.95 1.64 5.98 0.0459 16.9 1.88 31.77 0.1885 70.7 2.38 168.26
Approx. [12] 0.0058 3.89 1.62 6.30 0.0105 3.93 1.63 6.40 0.0445 16.8 1.87 31.42 0.1754 65.3 2.38 155.41
Approx. [6] 0.0056 3.60 1.54 5.54 0.0101 3.90 1.50 5.85 0.0432 15.8 1.86 29.39 0.1600 62.80 2.35 147.58
Approx. [5] 0.0057 3.80 1.44 5.47 0.0103 3.91 1.30 5.08 0.0440 16.2 1.80 29.16 0.1500 63.00 2.30 144.90
Proposed Approx. 0.0050 3.31 1.40 4.64 0.0090 3.79 1.27 4.82 0.0407 14.3 1.75 25.19 0.1312 53.92 2.23 120.26
8 Exact [6] 0.0191 6.38 3.36 21.44 0.0345 11.4 3.56 40.58 0.1363 49.8 3.61 179.78 0.5541 265.4 3.91 1037.71
Proposed Exact 0.0184 6.01 3.25 19.53 0.0333 11.0 3.42 37.62 0.1302 41.1 3.41 140.15 0.5498 242.6 3.91 948.56
Approx. [12] 0.0155 5.45 2.97 16.19 0.0301 10.4 3.31 34.42 0.1151 35.1 3.02 106.00 0.4424 193.7 3.88 751.556
Approx. [6] 0.0142 4.20 2.70 11.34 0.0290 9.60 2.90 27.84 0.1050 27.8 2.96 82.29 0.4200 166.0 3.70 614.20
Approx. [5] 0.0135 4.60 2.50 11.50 0.0285 9.20 2.55 23.46 0.1020 25.5 2.80 71.40 0.4000 150.0 3.40 510.00
Proposed Approx. 0.0108 3.45 2.42 8.36 0.0269 8.06 2.40 19.35 0.0955 20.5 2.74 56.18 0.3823 139.3 3.28 456.99
Figure 11: PADP comparison of existing designs [6, 5] and the proposed approximate design for different systolic array sizes
IV-B Error Analysis
Accuracy of the proposed approximate MAC design is evaluated using well-known error metrics [15] such as Error Distance (ED), Error Rate (ER), Relative Error Distance (RED), Mean Error Distance (MED), and Mean Relative Error Distance (MRED). These metrics measure the deviation of the approximate outputs from the exact. These metrics are estimated using uniformly distributed input combinations using Python simulations. Table V summarizes the error metrics of 4-bit and 8-bit PE designs. The results indicate that the accuracy of the designs gradually decreases with increasing approximation factor of $k$, although approximations up to $k=N$ bits still produce acceptable results as discussed in the next section.
TABLE V: Error Metrics for Proposed and Existing PE Designs
Column Wise Unsigned Signed
Approx NMED MRED NMED MRED
4 bit design
Prop Apx k=2 0.007 0.091 0.007 0.271
Prop Apx k=3 0.015 0.141 0.015 0.466
Prop Apx k=4 0.030 0.223 0.023 0.642
8 bit design
Prop Apx k=2 0.0001 0.0011 0.0001 0.0037
Prop Apx k=4 0.0004 0.0033 0.0004 0.0130
Prop Apx k=5 0.0006 0.0075 0.0006 0.0286
Prop Apx k=6 0.0018 0.0108 0.0018 0.0481
Prop Apx k=8 0.0077 0.0328 0.0073 0.2418
Design[6]* 0.00075 0.0082 0.00079 0.0364
Design[12]* 0.00072 0.0075 0.00086 0.0328
Design[5]* 0.00068 0.0068 0.00072 0.0292
  • *Five columns (k=5) are approximated.
As shown in Fig. 12, the proposed approximate design achieves a significant reduction in PDP while maintaining a lower MED compared to existing approximate designs. Although the design in [5] attains a slightly lower MED, the proposed design offers superior area, power, and delay metrics. Furthermore, the proposed exact design demonstrates improved efficiency up to 10% over the exact design  [6].
Figure 12: Comparison of PDP and MED for Signed 8-bit PE.
V Applications
In this section, we evaluate the effectiveness of the proposed approximate processing elements (PEs) through three representative applications. The image quality is estimated using metrics such as peak signal to noise ratio (PSNR) and structural similarity index measure (SSIM).
V-A Discrete Cosine Transform (DCT)
It is a widely used mathematical tool in image and video compression standards such as JPEG,and MPEG. In this paper, the proposed processing element (PE) of the systolic array is applied to perform matrix multiplication for an $\times8$ two-dimensional integer-scaled DCT. Instead of using floating-point operations, the DCT coefficient matrix is scaled to integer values [16], making it suitable for hardware architectures that operate on fixed-point or integer arithmetic. This approach reduces hardware complexity, power consumption, and area, while maintaining acceptable image reconstruction quality.
The input image, DCT coefficients, and reconstructed image are shown in Fig.  13. The Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM) are calculated between the original and reconstructed images. This simulation is implemented in a Python environment and employs an $\times8$ systolic array. The output images depicted in Fig. 13 (c) represents the (N/22)(N/2-2) column approximation. As the column-wise approximation parameter $k$ increases, the quality of the output images degrades in terms of PSNR and SSIM as shown in Table VI.
TABLE VI: Quality metrics of DCT for N=8 Processing Element
k Column Approx. PSNR (dB) SSIM
3 (N/21)(N/2-1) 38.21 0.955
4 $(N/2)$ 35.67 0.923
7 (N1)(N-1) 28.43 0.872
Refer to caption Refer to caption Refer to caption
(a) Input image (b) DCT coefficients (c) Reconstructed
(45.97 dB, 0.991)
Figure 13: Resultant images of DCT along with (PSNR and SSIM) values.
V-B Edge Detection
Edge detection is a fundamental step in image processing, widely used to extract object boundaries and structural information. Convolution-based operators, such as the Laplacian filter, are commonly employed for this task by emphasizing regions with high spatial intensity variation. In this work, we adopt the $3\times 3$ Laplacian kernel which is particularly effective for highlighting edges. However, the convolution operation requires multiple signed multiplications between pixel values and filter coefficients, making it computationally intensive.
To address this, the proposed approximate PE is integrated into the convolution process. The design supports different approximation column levels ($k=3,4,5,6$), where higher values of $k$ correspond to more aggressive approximations. As shown in Table VII, approximation levels of $k=3$ and $k=4$ achieve high PSNR values, indicating good image quality. Increasing the approximation level to $k=5$ and $k=6$ results in a slight degradation in image quality, but the edge structures remain well preserved while delivering greater hardware efficiency.
TABLE VII: Performance of Edge Detection using Approximate Signed Processing Element
Design Level of Approximation ($k$) PSNR (dB)
Proposed $k=2$ 35.45
$k=3$ 31.99
$k=4$ 30.12
$k=5$ 29.03
$k=6$ 28.41
Design [17] $k=3$ 31.78
Refer to caption Refer to caption Refer to caption Refer to caption
Input image Proposed exact k = 3 k = 4
PSNR: 31.99 dB PSNR: 30.12 dB
Figure 14: Edge detection results using the Laplacian filter with the proposed approximate signed PE at different approximation levels.
V-C Image Sharpening
Image sharpening enhances edges and fine details by applying convolution with a sharpening kernel. In this work, the proposed approximate PE is integrated to perform the convolution operation, enabling efficient sharpening while reducing computational cost compared to exact designs. Fig. 15 shows the output images obtained from the sharpening operation using the proposed approximate PE. The existing design reported in [17] achieves a PSNR of 27.12 dB.
Refer to caption Refer to caption Refer to caption Refer to caption
Input image Proposed exact k = 3 k = 4
(28.51 dB, 0.9804) (28.31 dB ,0.9634)
Figure 15: Image sharpening results using the sharpening filter with the proposed approximate signed PE at different approximation levels, along with (PSNR, SSIM).
VI Conclusion
This work presents an energy efficient systolic array architecture leveraging proposed exact and approximate PE with optimized PPC and NPPC blocks. The proposed designs demonstrate substantial energy savings of up to 32% while maintaining high output quality, validated through applications such as, DCT based image compression, convolution based edge detection and image sharping. These results highlight the suitability of the architecture for error-resilient image and vision processing tasks, offering a promising balance between performance, energy efficiency, and computational accuracy.