### v2

# Energy Efficient Exact and Approximate Systolic Array Architecture for Matrix Multiplication

Abstract
Deep Neural Networks (DNNs) require highly efficient matrix multiplication engines for complex computations. This paper presents a Systolic Array (SA) architecture incorporating novel exact and approximate Processing Elements (PEs), designed using energy-efficient partial product and nand based partial product cells, termed as PPC and NPPC, respectively. The proposed 8-bit exact and approximate PE designs are employed in a 8x8 SA, which achieves a energy savings of 16% and 68%, respectively, compared to the existing design. To demonstrate their effectiveness, the proposed signed PEs are integrated into a SA for Discrete Cosine Transform (DCT) computation, achieving high output quality with a PSNR of 45.97 dB. Furthermore, for edge detection applications using both kernel-based and CNN-based approaches, the proposed approximate signed PEs achieve PSNR values of 30.45 dB and 75.98 dB, respectively. These results highlight the potential of the proposed design to deliver significant energy efficiency while maintaining competitive output quality, making it well-suited for error-resilient image and vision processing applications.
I Introduction
Artificial Intelligence (AI) and Machine Learning (ML) have become integral to a wide range of applications, including computer vision, natural language processing, speech recognition, and autonomous systems [1, 2]. At the core of these workloads are deep neural networks (DNNs), which rely extensively on large-scale matrix multiplications to perform computations across multiple layers of interconnected neurons. Both the training and inference phases of DNNs involve billions of Multiply–ACcumulate (MAC) operations, making matrix multiplication one of the most computationally and energy-intensive kernels in modern AI accelerators [3].
Systolic arrays (SAs) have emerged as an effective hardware architecture for accelerating matrix multiplications [4]. Recent advances in domain-specific hardware, such as Google’s Tensor Processing Unit (TPU), have demonstrated the ability of systolic arrays to deliver high throughput for DNN workloads. However, conventional systolic arrays typically employ exact arithmetic units, which incur significant energy consumption and large silicon area overheads. These limitations hinder their deployment in resource-constrained platforms, such as edge devices and Internet of Things (IoT) systems, where energy efficiency is of paramount importance. To mitigate these challenges, approximate computing has been widely explored as a design paradigm that trades exactness for improvements in energy efficiency and hardware complexity.
In this paper, we propose both exact and approximate processing element that supports both signed and unsigned operations for systolic array-based matrix multiplication. The proposed PE significantly reduces hardware complexity and energy consumption while maintaining acceptable image quality. This makes it well-suited for error-tolerant applications such as image processing and machine learning, where efficiency is prioritized over exact precision. The major contributions of this paper are summarized as follows:
  • Proposed an energy-efficient approximate systolic array architecture for matrix multiplication, introducing novel exact and approximate Processing Elements (PEs).
  • Proposed approximate Partial Product Cell (PPC) and Nand based Partial Product Cell (NPPC), achieve 46.8% and 34.4% of energy savings respectively as compared to the best existing design [5].
  • Proposed exact signed 8-bit PE achieves 24.37% of energy saving over the design [6], while the proposed approximate signed 8-bit PE achieves 22.51% of energy saving over the design [5].
  • Validation of the proposed architecture on three applications: DCT for image compression, kernel-based edge detection and CNN-based edge detection, results have been found to achieve a PSNR of 45 dB for DCT, 30.45 dB for kernel-based edge detection and 75.98 dB for CNN-based edge detection.
The rest of the paper is organized as follows. Section II explains the matrix multiplication using systolic array architecture. Section III presents the proposed processing elements for the matrix multiplication. Sections IV and V present the results and applications, respectively. The conclusions are presented in Section VI.
II Systolic Array Based Matrix Multiplication Architecture
Matrix multiplication plays a major role in accelerating the applications such as machine learning, signal processing, and neural networks. Several systolic array architectures for matrix multiplication are found in literature [7, 8, 9]. Recently, with applications in neural networks with larger data, GEMM-based matrix multiplication using a systolic array is highly appreciated for machine learning applications [10], [8].
Figure 1: Architecture of $3\times 3$ Systolic Array for Matrix Multiplication[7].
The latency of the conventional systolic array architecture, illustrated in Fig. 1, is given by 3N23N-2 clock cycles [11].
In a conventional systolic array as shown in Fig. 1, the PE is the basic block for executing MAC operations [9]. Since a large number of PEs operate concurrently, it has a critical impact on the overall performance, energy efficiency, and area utilization. In the conventional design, a PE typically consists of a single multiplier followed by an adder to perform the MAC operation [9]. Alternatively, some studies optimize the PE by employing PPC and NPPC [6], to generate positive partial products and negative partial products, respectively, for enabling efficient signed multiplication as presented in Fig. 2. The partial products are then accumulated using full adders (FA) in the accumulation stage. Fig. 3 illustrates the exact PPC and NPPC blocks used to perform the bit-wise MAC operations.
Figure 2: Existing Design of Exact 4-bit Signed PE [6].
Figure 3: Conventional approach of Exact (a) PPC and (b) NPPC.
The next section presents the proposed processing elements for performing the matrix multiplication.
III Hardware Implementation of Proposed Processing Element (PE)
Firstly, we present the exact PE architecture using the proposed optimized PPC and NPPC cells then approximate PE architecture using approximate PPC and NPPC cells.
III-A Proposed Exact Processing Element
The proposed PE architecture eliminates the conventional separation between multiplication and accumulation by integrating them into a unified processing element to perform $a\times b+c$. This fused approach enables simultaneous reduction of both partial products and the accumulated sum, improving efficiency and reducing delay. The proposed exact PPC and NPPC for the PE are shown in Fig. 4 (a) and (b), respectively. Fig. 5 and 6 (b) illustrates the exact 8-bit and 4-bit signed PE architecture, respectively. The NPPC cells play a crucial role in handling signed operations.
Figure 4: Proposed exact designs of (a) PPC and (b) NPPC
Fig. 6 (a) illustrates the exact 4-bit unsigned PE. In this architecture, PPC play a crucial role in performing the MAC operation. For the 8-bit signed PE, a total of 14 NPPC cells and 50 PPC cells are required, whereas the existing work in [6] requires 15 number of additional full adders as well. In general, for an $N$-bit PE, the implementation requires $N^{2}-2N-2$ PPC cells and 2N22N-2 NPPC cells.
Figure 5: Proposed exact 8-bit signed PE Design.
Figure 6: Proposed 4-bit PE designs (a) exact unsigned, (b) exact signed, and (c) approximate signed (k=N1)(k=N-1) .
III-B Proposed Approximate Processing Element
In the proposed design, the PE is constructed using approximate PPC and NPPC blocks. Fig. 6 (c) illustrates the architecture of the proposed approximate 4-bit signed PE, where the dotted blocks represent the approximate PPC and NPPC units for approximate factor k=N1k=N-1. The proposed approximate PPC and NPPC are presented in Fig. 7, and their truth tables are provided in Table I.
The Boolean expression for the Sum output of the proposed approximate PPC cell is $S_{\text{out}}=\overline{\overline{(S_{\text{in}}+C_{\text{in}})}+(a_{i}\cdot b_{i})}$, while the corresponding carry output is $C_{\text{out}}=a_{i}\cdot b_{i}$. For the NPPC block, the Sum output is expressed as $S_{\text{out}}=\overline{(S_{\text{in}}+C_{\text{in})}\cdot\overline{(a_{i}\cdot b_{i})}}$, and the carry output is given by $C_{\text{out}}=(S_{\text{in}}+C_{\text{in})}\cdot\overline{(a_{i}\cdot b_{i})}$. The proposed approximate PPC introduces five erroneous outputs, out of a total of 16 cases, resulting in an error rate of $5/16$. The error occurs for the following combinations of inputs, $(a_{i},b_{i},S_{\text{in}},C_{\text{in}})$ are $(0,0,1,1)$, $(0,1,1,1)$, $(1,0,1,1)$, $(1,1,0,0)$, and $(1,1,1,1)$. The corresponding error probabilities ($P_{E}$) for these cases are $9/256$, $3/256$, $3/256$, $9/256$, and $1/256$, respectively. Therefore, the total error probability of the proposed PPC and NPPC blocks is $25/256$.
Figure 7: Proposed Approximate designs of (a) PPC and (b) NPPC.
TABLE I: Truth Table for Partial Product Cell (PPC) and Negative Partial Product Cell (NPPC)
Inputs PPC: $a\cdot b+C_{in}+S_{in}$ NPPC: $\overline{a\cdot b}+C_{in}+S_{in}$
Exact Approx Exact Approx
$a_{i}$ $b_{i}$ $C_{in}$ $S_{in}$ $C$ $S$ $C$ $S$ ED $C$ $S$ $C$ $S$ ED
0 0 0 0 0 0 0 0 0 0 1 0 1 0
0 0 0 1 0 1 0 1 0 1 0 1 0 0
0 0 1 0 0 1 0 1 0 1 0 1 0 0
0 0 1 1 1 0 0 1 -1 1 1 1 0 -1
0 1 0 0 0 0 0 0 0 0 1 0 1 0
0 1 0 1 0 1 0 1 0 1 0 1 0 0
0 1 1 0 0 1 0 1 0 1 0 1 0 0
0 1 1 1 1 0 0 1 -1 1 1 1 0 -1
1 0 0 0 0 0 0 0 0 0 1 0 1 0
1 0 0 1 0 1 0 1 0 1 0 1 0 0
1 0 1 0 0 1 0 1 0 1 0 1 0 0
1 0 1 1 1 0 0 1 -1 1 1 1 0 -1
1 1 0 0 0 1 1 0 +1 0 0 0 1 +1
1 1 0 1 1 0 1 0 0 0 1 0 1 0
1 1 1 0 1 0 1 0 0 0 1 0 1 0
1 1 1 1 1 1 1 0 -1 1 0 0 1 -1
IV Results and Discussion
This section presents hardware evaluation of the proposed and existing designs in terms of area, power, and delay using the Cadence Genus synthesis solution. In addition, error metrics are analyzed through Python-based simulations.
IV-A Hardware Evaluation
The existing and proposed designs are designed using Verilog HDL, and then the designs are synthesized using the Cadence Genus Synthesis solution using a 90-nm UMC technology. The synthesized results were then evaluated based on three primary hardware metrics: critical path delay, power consumption, and area, as summarized in Table II for the proposed PPC and NPPC. The analysis shows an energy improvement of 6.4% for the exact design [6], and a significant 46.8% improvement for the proposed approximate PPC design compared to the best existing design [12].
TABLE II: Hardware Analysis of Proposed and Existing PPCs and NPPCs
Design PPC NPPC
Area Power Delay PDP Area Power Delay PDP
[…]
Exact [6] 25.81 1.03 262 269.86 24.92 0.99 238 235.62
Prop Ext 24.98 0.99 255 252.45 23.47 0.99 216 213.84
Design [6] 13.32 0.64 187 119.04 12.54 0.61 156 95.16
Design [5] 14.13 0.58 157 91.06 13.22 0.60 148 88.80
Prop Apx 10.19 0.44 110 48.4 9.40 0.37 147 54.39
Table III presents the analysis of the proposed and existing PEs for both unsigned and signed designs. The results indicate that the proposed exact PEs achieve superior efficiency compared to existing exact designs, with notable reductions in area, power, and delay, leading to overall improvements of up to 16%. Table III also includes PE designs without PPC and NPPC blocks [13, 14]. A significant improvement in the Power Area Delay Product (PADP) of 65.45% is observed when compared to the design in [13], highlighting that PPC and NPPC based PE designs are highly efficient in terms of hardware metrics. Furthermore, the proposed approximate PE achieve up to 23% improvement in the PADP compared to the best existing approximate design [5].
TABLE III: Hardware Analysis of Proposed and Existing Processing Elements
PE bit Unsigned Signed
Design Width Area Power Delay PADPδ Area Power Delay PADPδ
[…]
Exact PPCs and NPPCs based Designs
Design 4 435.9 29.4 1.87 23.96 446.5 29.7 1.65 21.82
[6] 8 1718.5 181.3 3.92 1222.57 1708 183.4 3.71 1162.39
Design 4 432.8 30.4 1.76 23.13 445.3 31.7 1.55 21.88
[5] 8 1730.6 185.3 3.67 1175.71 1716 190.3 3.22 1050.21
Proposed 4 411 26.6 1.73 18.91 419 26.8 1.52 17.06
8 1659.2 180.7 3.65 1094.33 1620.3 170.6 3.18 879.02
Conventional Approach Exact MAC Designs
HA-FSA[10] 8 - - - - 2012 465 2.3 1662.1
Gemmini[13] 8 - - - - 1968 344 2.9 1763.7
Approximate PPCs and NPPCs based Designs (k=N1k=N-1)
Design [6] 4 416.30 24.1 1.56 15.64 435.9 29.6 1.69 21.78
8 1557.5 172.2 3.55 950.04 1546.3 216.0 3.51 1171.47
Design [12] 4 407.68 25.5 1.43 14.85 427.28 31.7 1.61 21.88
8 1476.2 164.1 3.21 777.51 1465.2 207.9 3.18 966.75
Design [5] 4 412.2 25.8 1.40 14.90 420.1 28.3 1.40 16.64
8 1012.1 145.5 3.01 442.91 975.5 177.2 2.50 431.93
Proposed 4 375.6 17.1 1.37 8.79 399.3 25.6 1.35 13.79
8 985.2 125.3 2.71 334.53 869.5 155.2 2.48 334.66
    : Reported values are normalized to 90 nm using DeepScale tool [15].
  • $\delta$
    : PADP unit is ($\mu\mathrm{m}^{2}\!\cdot\!\mathrm{fJ}$).
Table IV summarizes the results for various systolic array (SA) sizes at 250MHz250~MHz. Fig. 8 (a) shows up to 5.9% area savings and 14.1% energy savings of the proposed exact design over [6] across various sizes of systolic array. For all approximate SA configurations of varying sizes, each PE is used with an approximation factor of k=N1k=N-1. For the 8-bit configuration with a 16×16 SA, the proposed approximate design achieves a 62.7% reduction in PDP compared to the exact design [6] and a 24.2% improvement compared to the best existing approximate design [5] as shown in Fig. 8 (b). These results indicate that the proposed approximate designs achieve substantial gains in hardware efficiency, particularly for large systolic arrays, while maintaining competitive accuracy-performance tradeoffs.
Figure 8: Area and PDP comparison of the proposed design across different systolic array sizes using 8-bit signed PE, where grey line indicating the improvement in percentage for area and energy.
TABLE IV: Hardware Analysis of Different Signed PE Designs for Various Systolic Array Sizes at 250 MHz
bit Design 3×3 4×4 8×8 16×16
Width Area Power Delay PDP Area Power Delay PDP Area Power Delay PDP Area Power Delay PDP
(mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ) (mm2) (mW) (ns) (pJ)
4 Exact [6] 0.0062 3.98 1.65 6.57 0.0112 3.98 1.67 6.65 0.0465 17.2 1.88 32.34 0.1901 74.4 2.41 179.30
Proposed Exact 0.0060 3.90 1.63 6.35 0.0110 3.95 1.64 5.98 0.0459 16.9 1.88 31.77 0.1885 70.7 2.38 168.26
Approx. [12] 0.0058 3.89 1.62 6.30 0.0105 3.93 1.63 6.40 0.0445 16.8 1.87 31.42 0.1754 65.3 2.38 155.41
Approx. [6] 0.0056 3.60 1.54 5.54 0.0101 3.90 1.50 5.85 0.0432 15.8 1.86 29.39 0.1600 62.80 2.35 147.58
Approx. [5] 0.0057 3.80 1.44 5.47 0.0103 3.91 1.30 5.08 0.0440 16.2 1.80 29.16 0.1500 63.00 2.30 144.90
Proposed Approx. 0.0050 3.31 1.40 4.64 0.0090 3.79 1.27 4.82 0.0407 14.3 1.75 25.19 0.1312 53.92 2.23 120.26
8 Exact [6] 0.0191 6.38 3.36 21.44 0.0345 11.4 3.56 40.58 0.1363 49.8 3.61 179.78 0.5841 265.4 3.91 1037.71
Proposed Exact 0.0184 6.01 3.25 19.53 0.0333 11.0 3.42 37.62 0.1302 42.8 3.51 150.15 0.5498 233.3 3.82 891.3
Approx. [12] 0.0155 5.45 2.97 16.19 0.0301 10.4 3.31 34.42 0.1151 35.1 3.02 106.00 0.4424 193.7 3.88 751.556
Approx. [6] 0.0142 4.20 2.70 11.34 0.0290 9.60 2.90 27.84 0.1050 27.8 2.96 82.29 0.4200 166.0 3.70 614.20
Approx. [5] 0.0135 4.60 2.50 11.50 0.0285 9.20 2.55 23.46 0.1020 25.5 2.80 71.40 0.4000 150.0 3.40 510.00
Proposed Approx. 0.0110 3.86 2.42 9.36 0.0249 8.06 2.40 19.35 0.0895 20.5 2.74 56.18 0.3513 117.8 3.28 386.5
IV-B Error Analysis
Accuracy of the proposed approximate MAC design is evaluated using well-known error metrics [16], Mean Relative Error Distance (MRED) and Normalized Mean Error Distance (NMED). These metrics measure the deviation of the approximate outputs from the exact. These metrics are estimated using uniformly distributed input combinations for 8-bit PE (65536) using Python simulations. Table V summarizes the error metrics of 4-bit and 8-bit PE designs. The results indicate that the accuracy of the designs gradually decreases with increasing approximation factor of $k$, although approximations up to $k=N$ bits still produce acceptable results as discussed in the next section.
TABLE V: Error Metrics for Proposed and Existing 8-bit PE Designs
Design k Unsigned Signed
NMED MRED NMED MRED
Proposed 22 0.0001 0.0011 0.0001 0.0037
44 0.0004 0.0033 0.0004 0.0130
Approx. 55 0.0006 0.0075 0.0006 0.0286
66 0.0018 0.0108 0.0022 0.0481
88 0.0077 0.0328 0.0081 0.2418
Design[5] 66 0.0025 0.0146 0.0033 0.0626
Design [6] 66 0.0075 0.0824 0.0079 0.1064
Design[12] 66 0.0061 0.0250 0.0046 0.0758
As shown in Fig. 9, the proposed approximate design achieves a significant reduction in PDP while maintaining a lower NMED compared to existing approximate designs. Although the design in [5] attains a slightly lower NMED, the proposed design offers superior area, power, and delay metrics. Furthermore, Fig. 10 presents the variation of PDP and MRED values for different approximation levels for signed 8-bit PE.
Figure 9: Comparison of PDP and NMED for Signed 8-bit PE for k=N1k=N-1.
Figure 10: Variation of PDP and MRED with factor ($k$) for signed 8-bit PE.
V Applications
In this section, we evaluate the effectiveness of the proposed PEs through three representative applications. The first application implements an $\times8$ Discrete Cosine Transform (DCT) for image compression, while the second performs edge detection using a Laplacian kernel filter. The third application integrates the proposed approximate PEs into a Convolutional Neural Network (CNN) based Bi-Directional Cascade Network (BDCN) architecture [17] for edge detection. The image quality of the approximate PEs is evaluated using Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM), computed with respect to the outputs of the exact design.
V-A Discrete Cosine Transform (DCT)
DCT is widely used in image and video compression standards such as JPEG and MPEG. In this work, the proposed approximate PE in the systolic array performs matrix multiplication for an $\times8$ integer-scaled DCT. The DCT coefficient matrix is scaled to integer values [18], making it suitable for fixed-point hardware implementations, while effectively reducing computational complexity, power consumption, and area, without significantly compromising image quality.
The input, DCT coefficients, and reconstructed images are shown in Fig. 11 along with PSNR and SSIM value. Fig. 11 (c) shows the output for the $k=2$ column approximation using proposed 8-bit signed PE. As the approximation parameter $k$ increases, the output quality decreases in terms of PSNR and SSIM, as summarized in Table VI.
TABLE VI: Performance of approximate PE
PE k DCT Edge Detection BDCN-ED
Designs PSNR SSIM PSNR SSIM PSNR SSIM
Design [5] 8 27.97 0.854 10.97 0.538 28.56 0.865
Design [6] 8 26.15 0.792 9.12 0.424 17.29 0.798
Design [12] 8 26.48 0.837 9.24 0.471 20.29 0.819
Proposed 2 45.97 0.991 30.45 0.910 75.98 1
4 38.21 0.955 20.51 0.894 68.55 1
6 35.67 0.923 12.76 0.678 51.52 0.999
8 28.43 0.872 11.41 0.651 34.60 0.995
Refer to caption Refer to caption Refer to caption
(a) Input image (b) DCT coefficients (c) Reconstructed
(45.97 dB, 0.991)
Figure 11: Resultant images of DCT along with (PSNR and SSIM) values.
V-B Edge Detection
The proposed approximate signed PE is evaluated for edge detection using both conventional laplacian kernel and a CNN-based BDCN [17]. In the kernel-based approach, the laplacian filter is convolved using different approximation levels. Lower $k$ values provide high PSNR and SSIM with good visual quality, while higher $k$ values yield major degradation in the image quality as shown in Fig. 13 first row.
In the CNN-based approach, the proposed architecture is a modified version of the conventional Bi-Directional Cascade Network for perceptual edge detection (BDCN) model [17] as shown in Fig. 12. This network consists of a sequence of convolution, pooling, and upsampling layers arranged in a bi-directional cascade form. The first two blocks employ approximate PEs to reduce computation cost and power consumption, whereas the subsequent blocks maintain full-precision computations to preserve accuracy in the later feature extraction stages. This hybrid design effectively balances accuracy and efficiency, as the early layers can tolerate small computation errors.
Figure 12: Approximate Architecture of BDCN [17].
As observed from Table VI and Fig. 13, the CNN-based edge detection consistently achieves higher PSNR and SSIM than the kernel-based method. Even for higher approximation factors ($k=6,8$), BDCN maintains superior image quality, demonstrating that the network structure efficiently compensates for arithmetic inaccuracies while achieving significant power and area savings. This validates the effectiveness of integrating approximate PEs within CNN-based edge detection frameworks, highlighting their potential for energy-efficient deep learning hardware implementations.
Refer to caption Refer to caption Refer to caption Refer to caption
Input image Proposed exact Proposed $k=4$ Proposed $k=6$
using kernel PSNR: 20.51dB PSNR: 12.76 dB
Refer to caption Refer to caption Refer to caption Refer to caption
Proposed exact Proposed $k=2$ Proposed $k=8$ Design [12]
using BDCN PSNR: 75.98 dB PSNR: 34.60 dB PSNR: 20.29 dB
Figure 13: Comparison of edge detection results, where the first row presents the outputs generated by the Laplacian kernel, while the second row illustrates the corresponding results produced by the BDCN model.
VI Conclusion
This work presents an energy efficient systolic array architecture leveraging proposed exact and approximate PE with optimized PPC and NPPC blocks. The proposed designs demonstrate substantial energy savings of up to 68% while maintaining high output quality, validated through applications such as, DCT based image compression and convolution based edge detection. These results highlight the suitability of the architecture for error-resilient image and vision processing tasks, offering a promising balance between performance, energy efficiency, and computational accuracy.